DragonFlights, Inc. is planning on purchasing a new flying dragon for their new route to Volantis. The cost of the dragon isWhat are the risks related to the investment into new dragon?What is your managerial decision or recommendation about the new dragon project?
Answer #1

Cost of equity by using CAPM = Risk free rate+(Maret risk premium*Beta) = 2.5%+(7%*2.8) = 2.5%+19.6% = 22.1%

Cost of Redeemable Debenture (Kd) (RV-NP (1-t)+ n (RV+NP) 2 Where, - NP = Interest payment = Net proceeds from debentures in

Coupon(I) per 6months = Face value*Coupon rate/2 = 1000*14%/2 = 70

NP = 1,250 RV=1,000 t=0.25 n=20years*2 = 40 periods

Cost of debt after tax = {[70*(1-0.25)]+[(1,000-1,250)/40]}/[(1000+1250)/2] = {(70*0.75)+(-250/40)}/(2250/2) = (52.5-6.25)/1125 = 46.25/1125 = 4.11% per 6months period = 8.22% per annum

Computation of hurdle rate:

Type Market value Proportion (Market value/ΣMarket value) Cost after tax Cost of capital (Cost after tax *Proportion)
Common stock 116,000,000 65.00% 22.10% 14.37%
Debt 62,500,000 35.00% 8.22% 2.88%
178,500,000 17.24%

Cost of capital or hurdle rate = 17.24%

Sl No Years 0 1 2 3 4 5 6 7 8 9 10
i Revenue (Previous year*98%) 8,250,000 8,085,000 7,923,300 7,764,834 7,609,537 7,457,347 7,308,200 7,162,036 7,018,795 6,878,419
ii Annual Maintenance (Given) 150,000 150,000 150,000 150,000 150,000 150,000 150,000 150,000 150,000 150,000
iii Salary to pilot (Given) 250,000 250,000 250,000 250,000 250,000 250,000 250,000 250,000 250,000 250,000
iv Variable cost (Previous year*98%) 2,250,000 2,205,000 2,160,900 2,117,682 2,075,328 2,033,822 1,993,145 1,953,282 1,914,217 1,875,932
v Feeding cost (Given) 450,000 450,000 450,000 450,000 450,000 450,000 450,000 450,000 450,000 450,000
vi Other fixed cost (Given) 1,250,000 1,250,000 1,250,000 1,250,000 1,250,000 1,250,000 1,250,000 1,250,000 1,250,000 1,250,000
vii Depreciation (10.3million/10years) 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000
viii EBIT (i-ii-iii-iv-v-vi-vii) 2,870,000 2,750,000 2,632,400 2,517,152 2,404,209 2,293,525 2,185,054 2,078,753 1,974,578 1,872,487
ix Tax @ 25% (viii*0.25) 717,500 687,500 658,100 629,288 601,052 573,381 546,264 519,688 493,645 468,122
x PAT (viii-ix) 2,152,500 2,062,500 1,974,300 1,887,864 1,803,157 1,720,144 1,638,791 1,559,065 1,480,934 1,404,365
xi Add back: Depreciation 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000 1,030,000
xii Operating cash flow (x+xi) 3,182,500 3,092,500 3,004,300 2,917,864 2,833,157 2,750,144 2,668,791 2,589,065 2,510,934 2,434,365
xiii Investment (Given) -10,300,000
xiv Net cash flow (xii+xiii) -10,300,000 3,182,500 3,092,500 3,004,300 2,917,864 2,833,157 2,750,144 2,668,791 2,589,065 2,510,934 2,434,365
xv PVF @ 17.24% 1.00 0.8530 0.7276 0.6206 0.5293 0.4515 0.3851 0.3285 0.2802 0.2390 0.2039
xvi Net present value of cash flow (xiv*xv) -10,300,000 2,714,673 2,250,103 1,864,469 1,544,425 1,279,170 1,059,080 876,698 725,456 600,113 496,367

Net present value = ΣNet present value of cash flow = 3,110,554

Payback period:

Year Net cash flow

Cumumulative cash flow

0 -10,300,000 -10,300,000
1 3,182,500 -7,117,500
2 3,092,500 -4,025,000
3 3,004,300 -1,020,700
4 2,917,864 1,897,164
5 2,833,157 4,730,321

Payback period = Year before the year in which cummulative cashflow turns into positive+(1st positive cummulative cashflow/cashflow in the year in which cummulative cashflow turns into positive) = 3years+(1,897,164/2917,864) = 3years+0.65year = 3.65years

A B C D E F G H - J K L M 1 2 - 3 4 5 6 7 8 9 SI No Years Revenue Annual Maintenance jii Salary to pilot iv Variable cost V F

Management decision:
Since the Payback period is less than 4years, NPV is positive & IRR is more than hurdle rate hence accept the dragon project.